# Construction of Urban Greenland Resources Collaborative Management Platform

### v2

@@ 1 Introduction
Urban growth and rising living standards have made greening essential for improving the quality of life. In China, over 65% of the population lives in urban areas, yet the average park space per person is only 15.29 square meters, the recommended 20 square meters [1]. An effective greening management system is crucial for achieving "carbon peaking and neutrality." Access to green spaces can reduce depression risk by 20% and increase life satisfaction by 5.2 percentage points[2]. Access to green spaces is currently uneven; in cities like Shanghai, central areas have only 60% of the green space compared to suburban regions. Initiatives such as Shenzhen’s "Park City" project aim to address these gaps. Measuring the carbon Benefits of green spaces is vital, with Chengdu reporting a carbon sink value of 1.23 billion yuan. Programs like Beijing’s "Tree Adoption" involve citizens caring for green areas [3]. Innovative greening management systems present practical solutions to urban challenges by enhancing ecological capacity and improving the well-being of residents. This approach contributes to developing more livable cities, fostering a balanced coexistence between urban environments and natural ecosystems[4]. Foreign urban greening management systems in research and development and operation have accumulated some worthy of domestic urban greening management system research and development and reference results[5]. Various domestic urban greening management systems have been designed and developed utilizing different technologies and architectures. These systems typically offer user management, equipment management, and environmental data monitoring functionalities. However, significant room remains for improvement in system intelligence, analytical capabilities, and overall user experience[6]. Secondly, the domestic scientific research on happy urban greening management systems is still faced with the dilemma of fast technical updates and non-uniform standards, which require continuous improvement and upgrading of technology. Increase the technical research and development and standardization[7]. It adopts a frontend and backend separation architecture and realizes enterprise data’s comprehensive management and query. They also emphasize scalability and user experience, essential for the happy city green management system[8]. This will help improve user experience and efficiency. It realizes the intelligent management of academic affairs information. The design of system functions and architecture is studied [9].
@@ 2 System Analysis
The system is organized into three modules: The administrator, gardener, and supplier modules [10]. The administrator module is designed to configure user settings, manage equipment, and input information regarding the urban greening environment and plant data. It also oversees the management of purchase applications and conducts statistical analyses pertinent to urban greening efforts[11]. The gardener module allows users to set personal information, manage plant details, and perform maintenance operations. It also includes the functionality to review environmental data to develop appropriate maintenance strategies, as well as to carry out facility maintenance tasks. The supplier module is responsible for providing flowers and trees. Through effective supply management, this module ensures that resources are allocated efficiently to meet the requirements of customers [12].
@@ 3 System Architecture
This system uses a layered architecture design to build a modular, highly secure, scalable garden management platform[17]. The architecture is divided into five layers: The Web client layer is based on Vue.js to achieve a responsive front-end, adapting to the 4K resolution display requirements and providing differentiated interactive interfaces for multiple roles, such as administrators, gardeners, and vendors; the application layer relies on the Spring Boot framework to provide core business functions, such as user authentication, vegetation management (trees/flowers/green space), supply chain collaboration, and feedback processing, via RESTful API; and efficient front-end and back-end communication based on HTTP/HTTPS protocols[18]; the architecture is divided into five layers(See Figure4) .The application layer is developed based on the Spring Boot framework. RESTful APIs provide core business functions such as user authentication, vegetation management (trees/flowers/green areas), supply chain collaboration, and feedback processing[19]. It achieves efficient front-end and back-end communication based on HTTP/HTTPS protocols. The infrastructure layer has high-performance network equipment to support 4K content transmission and real-time business processing in high-concurrency scenarios[20].
@@
Functional design: Administrators can manage the entire supply chain, while gardeners are limited to vegetation maintenance operations[21]. The core business module encompasses the complete life cycle management of vegetation, including classification entry, status tracking, and data analysis[22]. It also facilitates closed-loop processing of user feedback, such as problem reporting and work order flow, while optimising the supplier collaboration process. Additionally, it integrates operation logging and security auditing functions to ensure business traceability[23]. In terms of non-functional requirements, the system strengthens security through HTTPS encrypted transmission, hash storage of sensitive information, and regular vulnerability scanning; it adopts Redis caching, load balancing, and microservices architecture to improve performance, supporting 1,000+ concurrent requests per second, with a response time of 500ms or less[24]; it achieves elasticity through the deployment of Docker containers and Kubernetes clusters, and ensures horizontal expansion[25]. Docker containerisation and Kubernetes cluster deployment enable flexible expansion and reduction of capacity, guaranteeing horizontal scalability; in addition, daily incremental backup of the database and a 30-minute fault recovery mechanism further enhance the system’s disaster recovery capability[26]..
@@
The HTTP protocol serves as the core communication hub in this system. As a standardized interaction protocol between the Web client (Vue.js front-end) and the application layer (Spring Boot back-end), it defines the request-response model[27]. It ensures decoupled collaboration between the front end and back end. It supports the realization of all business functions, including user authentication (login/registration), resource management (trees, flowers, and green spaces), feedback, and supply chain operations, all of which transmit commands and data via HTTP[28]. In terms of technical implementation, the front-end initiates HTTP requests (e.g., GET/POST) to the back-end service layer, which triggers the data layer (MySQL) to add, delete, change, and check operations and returns the results in JSON format. At the same time, HTTP integrates permission control (administrator, gardener, and vendor roles) and secure access through the cookie/Token mechanism. As a cornerstone of modern Web applications, the high compatibility and scalability of this protocol provide a stable and efficient communication foundation for the system[29].
@@
Regarding technical implementation, the system follows the principle of front-end and back-end separation, with Vue.js as the core of the front-end and Spring Boot framework in the back-end, and strictly aligns with OAuth2.0 and JWT standards to achieve authentication[30]. The deployment level relies on cloud-native technology to ensure cross-platform compatibility and ease of operation and maintenance[31]. In the end, through the output of complete architectural design documents, interface specifications, and operation and maintenance manuals, the system can efficiently support multi-role collaboration, complex business scenarios, and future functionality expansion needs with stability and flexibility[32].
@@ 3.1.1 Detailed Design of Database
Urban Greening Management System mainly contains seven core components: Administrator module, gardener module, supplier module, flower management, tree management, green space management, and purchase application module. Among them, the administrator module includes the administrator name, user name, password, and other information; the gardener personnel module consists of the gardener name, phone, password, and other information; the supplier module consists of the supplier company name, phone, password, and other information; flower management module includes flower name, photo description and other information; tree management module includes tree name, photo description and other information; purchase application module includes order number, Order details, shipping status and other information [33].
@@ 5 Evaluation
Testing can find the problem areas of the system. In the process of system development, to ensure the quality of the code, need to carry out various forms of testing of the system, which can help to find potential problems in the system, such as anomalies, loopholes, and so on [35]. Testing can guarantee delivery quality; testing can effectively improve the stability and reliability of the system to ensure that the system can be successfully delivered online and ensure good operation. Testing can improve user satisfaction. Testing can check the user’s perspective and further optimize their experience to enhance user satisfaction[36].
@@ 5.1 Use Cases
The system’s testing was focused on the following modules: Tree Management Module, Flower Management Module, Green Space Management Module, and Order Request Module. The testing covered functionality testing, which included adding, inserting, deleting, querying, and applying to ensure that each module worked properly under different functions[32]..
@@
Test case numbers are named after module functions and operation types, which clarify the test objectives. The accuracy and stability of the module functions are verified by testing different conditions and input data[37]. This test focuses on adding, inserting, deleting, and querying tree information in Tree Management, with test cases to show that the module functions as expected(See the TableLABEL:TC:1):

### v3

@@ 1 Introduction
Urban growth and rising living standards have made greening essential for improving the quality of life. In China, over 65% of the population lives in urban areas, yet the average park space per person is only 15.29 square meters, the recommended 20 square meters [49]. An effective greening management system is crucial for achieving "carbon peaking and neutrality." Access to green spaces can reduce depression risk by 20% and increase life satisfaction by 5.2 percentage points[16]. Access to green spaces is currently uneven; in cities like Shanghai, central areas have only 60% of the green space compared to suburban regions. Initiatives such as Shenzhen’s "Park City" project aim to address these gaps. Measuring the carbon Benefits of green spaces is vital, with Chengdu reporting a carbon sink value of 1.23 billion yuan. Programs like Beijing’s "Tree Adoption" involve citizens caring for green areas [64]. Innovative greening management systems present practical solutions to urban challenges by enhancing ecological capacity and improving the well-being of residents. This approach contributes to developing more livable cities, fostering a balanced coexistence between urban environments and natural ecosystems[34]. Foreign urban greening management systems in research and development and operation have accumulated some worthy of domestic urban greening management system research and development and reference results[31, 19]. Various domestic urban greening management systems have been designed and developed utilizing different technologies and architectures. These systems typically offer user management, equipment management, and environmental data monitoring functionalities. However, significant room remains for improvement in system intelligence, analytical capabilities, and overall user experience[52, 20]. Secondly, the domestic scientific research on happy urban greening management systems is still faced with the dilemma of fast technical updates and non-uniform standards, which require continuous improvement and upgrading of technology. Increase the technical research and development and standardization[53, 46]. It adopts a frontend and backend separation architecture and realizes enterprise data’s comprehensive management and query. They also emphasize scalability and user experience, essential for the happy city green management system[18, 62]. This will help improve user experience and efficiency. It realizes the intelligent management of academic affairs information. The design of system functions and architecture is studied [10].
@@ 2 System Analysis
The system is organized into three modules: The administrator, gardener, and supplier modules [50, 2]. The administrator module is designed to configure user settings, manage equipment, and input information regarding the urban greening environment and plant data. It also oversees the management of purchase applications and conducts statistical analyses pertinent to urban greening efforts[59, 7]. The gardener module allows users to set personal information, manage plant details, and perform maintenance operations. It also includes the functionality to review environmental data to develop appropriate maintenance strategies, as well as to carry out facility maintenance tasks. The supplier module is responsible for providing flowers and trees. Through effective supply management, this module ensures that resources are allocated efficiently to meet the requirements of customers [17].
@@ 3 System Architecture
This system uses a layered architecture design to build a modular, highly secure, scalable garden management platform[23, 51]. The architecture is divided into five layers: The Web client layer is based on Vue.js to achieve a responsive front-end, adapting to the 4K resolution display requirements and providing differentiated interactive interfaces for multiple roles, such as administrators, gardeners, and vendors; the application layer relies on the Spring Boot framework to provide core business functions, such as user authentication, vegetation management (trees/flowers/green space), supply chain collaboration, and feedback processing, via RESTful API; and efficient front-end and back-end communication based on HTTP/HTTPS protocols[54, 1]; the architecture is divided into five layers(See Figure4) .The application layer is developed based on the Spring Boot framework. RESTful APIs provide core business functions such as user authentication, vegetation management (trees/flowers/green areas), supply chain collaboration, and feedback processing[57, 42]. It achieves efficient front-end and back-end communication based on HTTP/HTTPS protocols. The infrastructure layer has high-performance network equipment to support 4K content transmission and real-time business processing in high-concurrency scenarios[60, 39].
@@
Functional design: Administrators can manage the entire supply chain, while gardeners are limited to vegetation maintenance operations[30, 12]. The core business module encompasses the complete life cycle management of vegetation, including classification entry, status tracking, and data analysis[63, 13]. It also facilitates closed-loop processing of user feedback, such as problem reporting and work order flow, while optimising the supplier collaboration process. Additionally, it integrates operation logging and security auditing functions to ensure business traceability[5, 15]. In terms of non-functional requirements, the system strengthens security through HTTPS encrypted transmission, hash storage of sensitive information, and regular vulnerability scanning; it adopts Redis caching, load balancing, and microservices architecture to improve performance, supporting 1,000+ concurrent requests per second, with a response time of 500ms or less[6, 22]; it achieves elasticity through the deployment of Docker containers and Kubernetes clusters, and ensures horizontal expansion[35, 38]. Docker containerisation and Kubernetes cluster deployment enable flexible expansion and reduction of capacity, guaranteeing horizontal scalability; in addition, daily incremental backup of the database and a 30-minute fault recovery mechanism further enhance the system’s disaster recovery capability[21, 43]..
@@
The HTTP protocol serves as the core communication hub in this system. As a standardized interaction protocol between the Web client (Vue.js front-end) and the application layer (Spring Boot back-end), it defines the request-response model[27, 45]. It ensures decoupled collaboration between the front end and back end. It supports the realization of all business functions, including user authentication (login/registration), resource management (trees, flowers, and green spaces), feedback, and supply chain operations, all of which transmit commands and data via HTTP[44, 55]. In terms of technical implementation, the front-end initiates HTTP requests (e.g., GET/POST) to the back-end service layer, which triggers the data layer (MySQL) to add, delete, change, and check operations and returns the results in JSON format. At the same time, HTTP integrates permission control (administrator, gardener, and vendor roles) and secure access through the cookie/Token mechanism. As a cornerstone of modern Web applications, the high compatibility and scalability of this protocol provide a stable and efficient communication foundation for the system[24, 4].
@@
Regarding technical implementation, the system follows the principle of front-end and back-end separation, with Vue.js as the core of the front-end and Spring Boot framework in the back-end, and strictly aligns with OAuth2.0 and JWT standards to achieve authentication[8, 14]. The deployment level relies on cloud-native technology to ensure cross-platform compatibility and ease of operation and maintenance[32, 47]. In the end, through the output of complete architectural design documents, interface specifications, and operation and maintenance manuals, the system can efficiently support multi-role collaboration, complex business scenarios, and future functionality expansion needs with stability and flexibility[25].
@@ 3.1.1 Detailed Design of Database
Urban Greening Management System mainly contains seven core components: Administrator module, gardener module, supplier module, flower management, tree management, green space management, and purchase application module. Among them, the administrator module includes the administrator name, user name, password, and other information; the gardener personnel module consists of the gardener name, phone, password, and other information; the supplier module consists of the supplier company name, phone, password, and other information; flower management module includes flower name, photo description and other information; tree management module includes tree name, photo description and other information; purchase application module includes order number, Order details, shipping status and other information [40, 3].
@@ 5 Evaluation
Testing can find the problem areas of the system. In the process of system development, to ensure the quality of the code, need to carry out various forms of testing of the system, which can help to find potential problems in the system, such as anomalies, loopholes, and so on [28, 11]. Testing can guarantee delivery quality; testing can effectively improve the stability and reliability of the system to ensure that the system can be successfully delivered online and ensure good operation. Testing can improve user satisfaction. Testing can check the user’s perspective and further optimize their experience to enhance user satisfaction[9, 26].
@@ 5.1 Use Cases
The system’s testing was focused on the following modules: Tree Management Module, Flower Management Module, Green Space Management Module, and Order Request Module. The testing covered functionality testing, which included adding, inserting, deleting, querying, and applying to ensure that each module worked properly under different functions[25, 36]..
@@
Test case numbers are named after module functions and operation types, which clarify the test objectives. The accuracy and stability of the module functions are verified by testing different conditions and input data[41, 48]. This test focuses on adding, inserting, deleting, and querying tree information in Tree Management, with test cases to show that the module functions as expected[56](See the TableLABEL:TC:1):